# DoraemonGPT : Toward Understanding Dynamic Scenes with Large Language Models (Exemplified as A Video Agent)

### v3

@@ 5 Conclusion
Regarding real-world scenarios’ dynamic and ever-changing nature, we present DoraemonGPT, an LLM-driven agent for solving dynamic video tasks. Compared to existing LLM-driven visual modular systems, DoraemonGPT has merits in: i) conceptually elegant system designed by delving into the dynamic modalities in our lives; ii) compact task-related symbolic memory by decoupling, extracting, and storing spatial-temporal attributes; iii) effective and decomposed memory querying through symbolic sub-task tools; iv) plug-and-play knowledge tools for accessing domain-specific knowledge; v) automated exploration of large planning space using MCTS planner, providing multiple solutions with an informative final answer; and vi) answer diversity that provides multiple potential candidates by fully exploring the solution space. Extensive experiments confirm the versatility and effectiveness of DoraemonGPT.
@@ Appendix A Appendix
§A.1 offers more implementation details of the MCTS planner.
§A.2 introduces several in-the-wild examples.
§A.5 discusses used foundation models.
§A.6 analyzes time of inference and efficiency of token usage.
§A.7 discusses our limitations.
§A.8 discusses the broader impacts of our work.
@@ A.2 In-the-wild Examples
In Fig. 7(a), we visualize the reasoning path of a standard video understanding task. As depicted, DoraemonGPT is asked to identify the speaker and analyze information about the dismissal. After several calls to various tools, DoraemonGPT got the right answers. Here we also visualize the time-dominant symbolic memory, which is the pivotal part of data processing in DoraemonGPT. Combining it with the well-defined symbolic language (SQL) promises transparency and efficiency.
@@ A.3 Inference Results on NExT-QA
A.4 Qualitative Results on Ref-YouTube-VOS
We further compared DoraemonGPT with ViperGPT in referring object segmentation on Ref-YouTube-VOS (Seo et al., 2020). As depicted in Fig. 9, our DoraemonGPT demonstrates a higher level of accuracy in identifying, tracking, and segmenting the referred objects mentioned by the users. In contrast, ViperGPT encounters failure cases where the recognized objects do not completely match the semantic and descriptive aspects. For example, when segmenting a person riding a pony, ViperGPT might inadvertently include the pony in the segmentation. Similarly, when segmenting a person wearing a hat, ViperGPT might overlook the region corresponding to the hat. These observations further validate the effectiveness of task-related symbolic memory in our DoraemonGPT model.
A.5 Discussion on the Impact of Foundation Models
@@
A.5.1 In space-dominant memory
@@
Referring Object Detection (Grounding DINO (Liu et al., 2023b)): DoraemonGPT harnesses the power of Grounding Dino to detect objects regarding their textual descriptions. For converting the detected bounding boxes into segmentation masks, SAM (Kirillov et al., 2023) is used. However, when facing complex user descriptions that involve multiple features (such as color, location, and size combined) or scenarios with multiple objects having similar semantics, Grounding Dino may encounter challenges in accurately detecting the desired object.
@@
A.5.2 In time-dominant memory
Speech recognition (Whisper (Radford et al., 2023)): Current methods can accurately convert audio to text. However, in multi-party conversation scenarios, the methods still cannot accurately perform voiceprint recognition for multiple speakers and accurately separate the results of different speakers. Additionally, it is challenging to match multiple voiceprints with the visual IDs of the speakers. This limitation restricts the ability of DoraemonGPT to infer and deduce the identities of speakers in complex multi-party conversation scenarios, relying solely on the inherent capabilities of LLMs.
Optical character recognition (OCR (PaddlePaddle, 2023) ): OCR technology can accurately recognize subtitles and well-structured text. However, it still struggles to robustly handle occluded text and artistic fonts.
@@ A.5.2 In time-dominant memory
A.6 Evaluation on the Inference Time and Token Usage Efficiency
@@
Due to the influence of simultaneous requests and network delay on ChatGPT’s online server, it’s impossible to fairly record the run-time of ChatGPT. Thus, a more equitable efficiency comparison when calling ChatGPT is to record the number of tokens used. As shown in the table above, DoraemonGPT’s prompt design is more efficient (617 tokens), which is less than VideoChat’s approach of directly incorporating video memory into the prompt (722 tokens) and significantly less than ViperGPT’s approach of including a large code definition in the prompt (4127 tokens). Additionally, even though the introduction of our MCTS planner divides the task into multiple nodes/steps, DoraemonGPT still requires far fewer tokens on average to obtain an answer compared to ViperGPT (1498 tokens vs 4127 tokens). Furthermore, DoraemonGPT significantly outperforms VideoChat (55.7 vs 51.8) on the challenging NExT-QA dataset.
@@ A.7 Limitations
Despite its comprehensive and conceptually elegant system, DoraemonGPT has some limitations for future studies. First, although TSM is a simple and effective way to decouple and handle spatial-temporal reasoning and DoraemonGPT has shown effectiveness with two task-related memory types (space-dominant and time-dominant), we believe that by further subdividing the types of tasks, we can introduce more nuanced categories of memory (e.g., human-centric memory) to construct task-related information with greater task-relevance. However, at present, the design of memory types is still a heuristic and manually driven process, lacking an automated design method. Second, the establishment of memory relies on the available foundation models (e.g., BLIP-2 (Li et al., 2023a)). In other words, foundation models’ performance directly influences memory’s reliability. Incorrect model predictions will introduce noise into the memory, thereby reducing its reliability and affecting the accuracy of decision-making. Additionally, foundation models may struggle to effectively extract the required video attributes in real-world scenarios that are difficult to generalize (e.g., low light, blurriness, occlusions, etc.). Third, the accuracy of planning in DoraemonGPT is limited by the capabilities of LLMs. When using a small-scale or insufficiently trained LLM, the likelihood of DoraemonGPT exploring reasonable solutions may be significantly reduced. Last, while the MCTS planner significantly improves the decision making ability of DoraemonGPT, it also introduces additional computational cost. This means that DoraemonGPT may only be available on high-end computing systems or online LLM services (OpenAI, 2021), limiting its use in real-time, resource-constrained scenarios.
A.8 Broader Impacts
DoraemonGPT aims to solve real-world dynamic tasks with LLMs and can handle video-based reasoning tasks, potentially revolutionizing several fields. Our system has potential applications in autonomous vehicles, surveillance systems, and interactive robotics, where dynamic understanding and decision making are crucial. However, it is important to consider the ethical implications and potential misuse of such systems. First, like many AI systems, DoraemonGPT could be exploited by malicious individuals for video manipulation or generating misleading content, posing threats to privacy and security. Protecting against such potential misuse requires robust safeguards and measures to detect and prevent malicious activities. Second, biases in the training data of LLMs or foundation models could unintentionally perpetuate discriminatory behavior. Mitigating biases and promoting fairness in the training and deployment of DoraemonGPT is essential to ensure equitable outcomes. Third, the reliance on external knowledge sources highlights the importance of data access and usage rights. Users and developers must adhere to regulations and ethical guidelines associated with these resources to avoid any legal complications. Fourth, the methodology introduced in DoraemonGPT holds potential for application of LLM-driven agents beyond the realm of vision. The rapid expansion of LLM-driven agents opens doors to transformative impacts across various fields (Wang et al., 2023c; Yuan et al., 2023; Park et al., 2023; Ruan et al., 2023; Qin et al., 2023; Shinn et al., 2023; Xu et al., 2023). DoraemonGPT, with its novel approach to modeling the dynamic aspects of visual scenes, tackles complex tasks through a computer vision lens. This innovation could extend its influence to other domains. For instance, in tool usage, our MCTS planner can offer effective exploration strategies in large solution spaces. Additionally, when it comes to open-world environments, our symbolic memory could provide precise guidances through symbolic language. This is particularly relevant for interactive planning scenarios(Park et al., 2023; Wang et al., 2023c).

### v4

@@ 5 Conclusion
Regarding real-world scenarios’ dynamic and ever-changing nature, we present DoraemonGPT, an LLM-driven agent for solving dynamic video tasks. Compared to existing LLM-driven visual modular systems, DoraemonGPT has merits in: i) conceptually elegant system designed by delving into the dynamic modalities in our lives; ii) compact task-related symbolic memory by decoupling, extracting, and storing spatial-temporal attributes; iii) effective and decomposed memory querying through symbolic sub-task tools; iv) plug-and-play knowledge tools for accessing domain-specific knowledge; v) automated exploration of large planning space using the MCTS planner, providing multiple solutions with an informative final answer; and vi) answer diversity that provides multiple potential candidates by fully exploring the solution space. Extensive experiments confirm the versatility and effectiveness of DoraemonGPT.
Acknowledgments
This work was supported by the National Science and Technology Major Project (No. 2023ZD0121300), the National Natural Science Foundation of China (No. 62372405), and the CCF-Tencent Open Fund.
Impact Statement
DoraemonGPT leverages LLMs to address real-world dynamic tasks, excelling in video-based reasoning with potential applications in autonomous vehicles, surveillance, and interactive robotics. Despite its promise, several ethical concerns must be addressed. i) The system could be misused for video manipulation or generating misleading content, necessitating robust safeguards against malicious activities. ii) Biases in training data could perpetuate discriminatory behavior, highlighting the need for fairness and bias mitigation. iii) Reliance on external knowledge sources underscores the importance of adhering to data access and usage regulations to avoid legal issues. iv) DoraemonGPT’s methodology extends the application of LLM-driven agents beyond vision, offering transformative impacts in various fields. For instance, the MCTS planner enhances exploration strategies in large solution spaces, while its symbolic memory aids interactive planning scenarios with precise guidance, relevant for diverse applications, e.g., embodied intelligence (Liu et al., 2024, 2023b; Dong et al., 2024).
@@ A.1 Implementation Details of MCTS Planner
The construction of a task-related knowledge base is also based on in-context learning. Concretely, we provide DoraemonGPT with in-context descriptions, usage, and examples of the two types of task-related symbolic memory, supplemented by the input question. The prompt is similar to Fig. 6, but the {tool_descriptions} is replaced with a {memory_descriptions}. As such, DoraemonGPT can autonomously construct one or both of the task-related knowledge bases. The attributes in each task-related symbolic memory are manually designed.
@@ A.2 In-the-wild Examples
In Fig. 8(a), we visualize the reasoning path of a standard video understanding task. As depicted, DoraemonGPT is asked to identify the speaker and analyze information about the dismissal. After several calls to various tools, DoraemonGPT got the right answers. Here, we also visualize the time-dominant symbolic memory, which is the pivotal part of data processing in DoraemonGPT. Combining it with the well-defined symbolic language (SQL) promises transparency and efficiency.
@@ A.3 Inference Results on NExT-QA
A.4 Discussion on the Impact of Foundation Models
@@
A.4.1 In space-dominant memory
@@
Referring Object Detection (Grounding DINO (Liu et al., 2023c)): DoraemonGPT harnesses the power of Grounding Dino to detect objects regarding their textual descriptions. For converting the detected bounding boxes into segmentation masks, SAM (Kirillov et al., 2023) is used. However, when facing complex user descriptions that involve multiple features (such as color, location, and size combined) or scenarios with multiple objects having similar semantics, Grounding Dino may encounter challenges in accurately detecting the desired object.
@@
A.4.2 In time-dominant memory
Speech recognition (Whisper (Radford et al., 2023)): Current methods can accurately convert audio to text. However, in multi-party conversation scenarios, the methods still cannot accurately perform voiceprint recognition for multiple speakers and accurately separate the results of different speakers. Additionally, it is challenging to match multiple voice prints with the visual IDs of the speakers. This limitation restricts DoraemonGPT’s ability to infer and deduce speakers’ identities in complex multi-party conversation scenarios, relying solely on the inherent capabilities of LLMs.
Optical character recognition (OCR (PaddlePaddle, 2023) ): OCR technology can accurately recognize subtitles and well-structured text. However, it still struggles to handle occluded text and artistic fonts robustly.
@@ A.4.2 In time-dominant memory
A.5 Evaluation on the Inference Time and Token Usage Efficiency
@@
Due to the influence of simultaneous requests and network delay on ChatGPT’s online server, it’s impossible to record ChatGPT’s run-time fairly. Thus, a more equitable efficiency comparison when calling ChatGPT is to record the number of tokens used. As shown in the table above, DoraemonGPT’s prompt design is more efficient (617 tokens), which is less than VideoChat’s approach of directly incorporating video memory into the prompt (722 tokens) and significantly less than ViperGPT’s approach of including a large code definition in the prompt (4127 tokens). Additionally, even though the introduction of our MCTS planner divides the task into multiple nodes/steps, DoraemonGPT still requires far fewer tokens on average to obtain an answer compared to ViperGPT (1498 tokens vs 4127 tokens). Furthermore, DoraemonGPT significantly outperforms VideoChat (55.7 vs 51.8) on the challenging NExT-QA dataset.
A.6 Typical Failure Cases
Here we list two typical failure cases: i) The foundation model cannot extract useful information for reasoning (see the top half of Fig. 10). ii) The correct answer is identified in the early exploration and the subsequent exploration results in wrong answers (see the bottom half of Fig. 10).
@@ A.7 Limitations
Despite its comprehensive and conceptually elegant system, DoraemonGPT has some limitations for future studies. First, although TSM is a simple and effective way to decouple and handle spatial-temporal reasoning and DoraemonGPT has shown effectiveness with two task-related memory types (space-dominant and time-dominant), we believe that by further subdividing the types of tasks, we can introduce more nuanced categories of memory (e.g., human-centric memory) to construct task-related information with greater task-relevance. However, at present, the design of memory types is still a heuristic and manually driven process, lacking an automated design method. Second, the establishment of memory relies on the available foundation models (e.g., BLIP-2 (Li et al., 2023a)). In other words, the performance of foundation models directly influences memory reliability. Incorrect model predictions will introduce noise into the memory, thereby reducing its reliability and affecting the accuracy of decision-making. Additionally, foundation models may struggle to effectively extract the required video attributes in real-world scenarios that are difficult to generalize (e.g., low light, blurriness, occlusions, etc.). Third, the accuracy of planning in DoraemonGPT is limited by the capabilities of LLMs. When using a small-scale or insufficiently trained LLM, the likelihood of DoraemonGPT exploring reasonable solutions may be significantly reduced. Last, while the MCTS planner significantly improves the decision-making ability of DoraemonGPT, it also introduces additional computational costs. This means that DoraemonGPT may only be available on high-end computing systems or online LLM services (OpenAI, 2021), limiting its use in real-time, resource-constrained scenarios.